# Richard A Berk

### v2

@@ 5 Surface Temperature Changes from Pre-Test to Post-Test
We transition to empirical analyses by separately examining the PNW heat wave summer temperatures and routine PNW summer temperatures. The response variable is gain scores of means grid cell by grid cell. For each grid cell, the gain score is the post-test mean temperature minus the pre-test mean temperature. It is just a usual difference in means that is computed separately for the actual heat wave and for each faux heat wave. Despite the name, gain scores can be positive or negatives and are sometimes called “change scores;” a gain score is nothing more than the change from pre-test to post-test. For the analyses immediately to follow, attention centers on the change for each grid cell in the mean surface temperature between the 4 day pre-test and the mean surface temperature during the 4 day post-test.555 Gain scores do not have to be a difference in means. They are more commonly a difference in scores, one for a pre-test and one for a post-test. For example, using a single temperature value the day before a heat wave and another single temperature value the day that heat wave begins would be an example. The length of their temporal separation would depend on the application.
Some grid cells are routinely warmer than others because of fix features such as urban heat islands. Gain scores control for fixed pre-test temperature differences between grid cells that might otherwise might be folded into post-test temperatures. Moreover, by removing earlier surface temperature variation that might be properly characterized as a nuisance in the study of heat waves, more accurate forecasts can result. These are a well-known properties of gain scores (Maris, 1998, Maxwell et al., 2018)
@@
5.1 Gain Score Distributions for the Actual Heat Wave and Faux Heat Wave 1
@@
Figure 3 shows the grid cell, surface temperature gain scores of means histogram for the extreme heat wave, for the Faux Heat Wave 1 and for their overlap. Variation in both sets of gain scores over grid cells is to be expected. The gain scores condition directly on pre-test surface temperatures and indirectly on fixed temperature-related features of grid cells such that dynamic forcings influencing surface temperatures can be better revealed.666 Statistical interactions between dynamic processes and fixed features of a grid cell certainly can occur and are not controlled by the pre-test conditioning. Estimating such relationships can be very challenging and is beyond the goals of this analysis, necessarily constrained by what the AIRS data measures.
@@
Neither the actual heat wave histogram nor faux heat wave histogram has much resemblance to the canonical normal distribution. There is far too much mass toward the centers of both histograms. Nor is either distribution strongly skewed to the left or the right. There is nothing to suggest an extreme value distribution. The attempt to isolate the role of dynamic temperature process does not lead to popular “off-the shelf” distributions in gain scores.777 These are “eyeball” assessments that require some care because the aspect ratio of all three histograms produces rather flat-looking frequency profiles. When the histograms have an aspect ratio of 1 to 1 in height to width, the eyeball impression is close to a rectangle that is taller than it is wide. The 1 to 1 aspect ratio led to distorted displays when collected on a single page.
One might conclude that the two histograms are much alike. However, the extreme heat wave histogram has 3 of 120 grid cells with negative values and a gain score mean of 3.8. The faux heat wave distribution has 44 of 128 grid cells with negative values and a gain score means of 1.2.888 There are less than 144 grid cells in either histogram because when a dataset for each was built with all of the predictors, there are 20 or more observations with missing data. One can explore further how the distributions compare by plotting the two histograms in the same figure. The bottom display shows the distributional offset in more detail. The red bars are taken from the histogram for the extreme heat wave gain scores. The blue bars are taken from the histogram for the faux heat wave 1 gain scores. The purple bars show the overlap. Clearly, the two distributions are quite different.
@@
In the interest of space, Faux Heat Wave 2 is not yet considered. But Its histogram looks a lot like the histogram for Faux Heat Wave 1. We make more use of Faux Heat Wave 2 later to replicate some important findings for Faux Heat Wave 1. The key point now is that when examining histograms, the role of a 4 day lag versus a 30 day lead does not seem to make much empirical difference in the gain score histograms.
6 Fitting The Gain Scores of Means With AIRS Predictors
@@
Differences between the two sets surface temperature gain scores have been described. As a step toward forecasting heat waves, one might wonder if the actual heat wave and Faux Heat Wave 1 have precursors that differ as well. Those precursors were shown in Table reftab: variables. To proceed, however, we need first to consider a bit further how the rare events problem might be mitigated.
@@
6.1 Mitigating The Rare Events Problem
@@
We have emphasized that extreme heat waves are statistically rare. But how rare depends on the population of which they are a part. For example, if one considers all days from 2017 through 2021 in the Pacific Northwest as a population of interest, there would be one extreme heat wave. Labeling all four of the last days in June of 2021 as heat wave days, there would be 4 extreme heat wave days for more than 1800 days overall (i.e., $365\times 5$). If one never predicted an extreme heat wave over all of those days, that prediction would be correct about 99.8% of the time without using any predictors at all. It would be, therefore, almost impossible to uncover the roles of predictors because fitting is already close to perfect. As noted earlier, one is stuck with very accurate and very trivial forecasts. Even if the data were limited to all of the days in 2021, a forecast of no heat wave would be correct 98.9% of the time. The fitting problems are nearly as catastrophic.
@@
There is more promise using data only for the 30 days in June of 2021. The trivial forecast then will be right “only” 86.7% of the time. We can make further progress by using grid cells as the study units for the post-test, pre-test, and on the day two weeks earlier when the predictors are measured. Figure 2 shows the result for the true heat wave, Faux Heat Wave 1, and Faux Heat Wave 2.
@@
The point of Figure 2 is solve the rare event problem across the three settings. If the actual heat wave scenario is compared to either of the two faux heat waves, there are in principle 144 true heat wave observations and 144 faux heat wave observations. If the extreme heat wave setting is compared to a pooled version of the two faux heat waves, there remain theoretically 144 observations in the heat wave scenario and then, 288 observations in the combined faux heat wave scenarios. In either case, the PNW extreme heat wave is not rare.
We are capitalizing on a variant of crossover, quasi-experimental designs (Lui, 2016; Maxwell et al., 2017). Crossover designs are common, especially in biomedical research and academic psychology. From Figure 2, all grid cells are exposed to extreme heat wave precursors at some specified time and the same grid cells are exposed to routine summer conditions at some other specified time. Because the grid cells serve as their own controls, subsequent forecasting accuracy is likely to improve.
@@
However, there is a price to pay for these design adjustments to the rare events problem. Perhaps most important, we use a form endogenous sampling when the three design configurations in Figure 2 are imposed to all grid cells. It is unlikely that any study of rare extreme heat waves could be undertaken in a naturally-occurring population such that one-third of the data were found in each of three different study scenarios. Forcing the data to be balanced in this manner can produce misleading statistical results (Manski and Lerman, 1977; Manski and McFadden, 1981). We will return to this problem later to consider several possible solutions. For now we continue to descriptively unpack how extreme heat wave surface temperatures can differ from faux heat wave surface temperatures.
6.2 An Application of Supervised Learning
We deploy a supervised learning algorithm that can have good forecasting skill. Just as shown in Figure 2, the response variable is grid cell gain scores of means for the true PNW heat wave and Faux Heat Wave 1. (Faux Heat Wave 2 will be used later.) The predictor values are taken from two weeks earlier for all variables shown in Table 1 except ”SurfAirTemp”, which is used to construct the pre-test and post-test. There are 286 observations, 144 grid cells each from the extreme and faux heat wave less two grid cells having missing data.999 Two grid cells were deleted because of missing data. We considered two forms of widely used and successful supervised learning: random forests (Breiman, 2001a) and stochastic gradient boosting (Friedman, 2001). Both can be seen as variants of nonparametric regression, are very easy to train, and often the default tuning parameter settings work well. A complication is that there is strong dependence between many of the AIRS predictors. For example, the measured temperatures at different altitudes are often correlated at values above .95.
Random forests handles strongly correlated predictors better than stochastic gradient boosting because it is comprised of a large ensemble of regression or classification trees with the partitioning for each tree undertaken using a small random subsets of predictors. Sampling predictors (without replacement) for each split can dramatically reduce the impact of high multicollinearity. We settled on the implementation of random forests in R, ported by Andy Law and Matthew Wiener (Liaw and Weiner, 2002) from the original Fortran by Breiman and Cutler (Breiman, 2001a), which includes several important auxiliary algorithms that can dramatically increase algorithmic transparency.101010 We briefly considered using deep neural networks. But the vastly different units for some of the predictors caused problems for the back propagation. Dependencies between the predictors also caused computational difficulties. Once both problems were addressed, and tuning was able to proceed, the fitting performance was a bit worse than for random forests and gradient boosting. Also many of the auxiliary procedures that shine some light into the algorithm black box (Molnar, 2022) were no longer readily available.
In this highly exploratory phase of the work, we were motivated by possible differences between how the same predictors performed depending on whether the data came from the true extreme heat wave or a faux heat wave. Work on quasiresonant amplification is a provocative illustration (Petoukov et al., 2012; Mann et al., 2018; Li et al., 2024). In response, we initially applied random forests separately to the extreme heat wave dataset and to the faux heat wave dataset structured as in faux heat wave 1. We computed 500 regression trees for each. About 75% of the gain score variance was accounted for using either the true heat wave data or the faux heat wave data. By the MSE criterion, fitting performance was promising and much the same for both.
But all else was quite different. For example, the fitted values for the heat wave histogram were farther to the right than the fitted values histogram for the faux heat wave, which follows in part from the different gain score distributions discussed earlier. Perhaps most telling, different predictors dominated the two different fits depending on whether the response variable was the gain scores for the extreme heat wave or the gain scores for the faux heat wave.
Figure 4: Variable importance plot for the random forest fit of the extreme heat wave (top) and Faux Heat Wave 1 (bottom). For both, the five more important proctors are shown. The variable names were shortened for ease of writing code but are more fully explained in the text.
Figure 4 is the variable importance plot for the extreme heat wave on the top and Faux Heat Wave 1 on the bottom. The five most important predictors for each are shown. The 14 day lagged predictors are listed in the left margin. Predictor importance is greater for predictors higher in listing. The metric at the bottom is the percentage increase in the gain score mean squared error (MSE) when each variable by itself is randomly shuffled such that associations with the gain scores and with the other predictors are driven toward zero. The increase in MSE measures the decrement in fitting performance.
Starting with the top figure at the top of variable list, when a grid cell’s topography (meters above sea level) is shuffled, the MSE increases by about 3%. If a grid cell’s land/sea indicator is shuffled, the MSE increases by about 1.7%. If two temperatures at relatively high altitudes are shuffled, MSE increases about 1.5% (i.e., 1% + .5%). If water vapor mass-mixing ratios in a middle altitude is shuffled, the MSE increases by about .5%. Although these are for the five largest increases in fitting error, in numerical terms they are quite small relative to the overall variance accounted for by the random forest. A significant reason is that the predictor contributions are shared with other predictors due to interaction effects that are not included in these MSE increases. This explains why the fixed grid cell features of elevation above sea level and whether grid cell is over land or over water appear in the figure.111111 It is possible to show these effects (Molnar, 2022), but the number of predictor pairs, triplets, and more that would need to be examined make this strategy computationally impractical. Interactions between predictors are included in the overall fit even though we cannot separately identify and characterize them.
The bottom display in Figure 4 is the variable importance plot for the faux heat wave. The two fixed predictors not shown because they have very small contributions to the fit. What matters most are several measures of temperature at different altitudes 14 days earlier. Grid cells that have larger gain scores during the extreme heat wave are warmer 14 days earlier as well. There is some form of temperature temporal dependence. But the point is that despite comparable fitting performance, the five predictors that uniquely matter most for the fit appear to differ depending on whether random forests is applied to the true heat wave data or the faux heat wave data.
In summary, both variable importance plots show which predictors are driving most the random forests’ fitted values. The predominant predictors differ depending on whether the gains scores belong to the actual heat wave or to the faux heat wave. This also is consistent with the claim that the extreme summer heat is not just an unusually harsh version of ordinary summer temperatures.
@@ 6.2 An Application of Supervised Learning
However, there is substantial dependence among the AIRS predictors. We repeated the random forest fitting several time for both the extreme heat wave and faux heat wave. The modest amount of randomness built into random forests sometimes led to disconcerting changes in the most important predictors identified. This instability motivated a look at faux heat wave 2. One can apply the exact same design used earlier but replace faux heat wave 1 with faux heat wave 2 and apply random forest as before. The crossover design now separates the actual and faux heat wave by 30 days instead of 4 days. Consequently the risks from spillover contamination are dramatically reduced. But the large gap in time can allow for differences in the grid cell data making comparisons between the predictors of the actual and faux heat waves more complicated. For example, on the average, weather conditions in June and July related to surface temperatures in may differ. There can be less rain and lower soil moisture in July, for instance. According to NOAA National Centers of Environmental Information, 1.9 inches of rain fell in Seattle in June of 2021, and 0.0 inches of rain fell in Seattle in July of 2021 (https://www.ngdc.noaa.gov/).121212 In the winter months, the rainfall in Seattle is much more substantial.
Despite such possibilities, the random forest fit to the faux heat wave gain scores of means in July is again quite good. About 70% of the variance in means-based gain scores was accounted for. Variable importance calculations determined the top five predictors as temperatures at middle to high altitudes, just as for the earlier faux wave. No water vapor mixing ratio made the cut. These results are similar to those for faux heat wave 1, despite some differences in the research design. Perhaps neither the difference in pre-test to post-test separation time nor the differences in weather patterns between June and July matter much. And there is some additional evidence that extreme summer heat waves may differ structurally for ordinary summer heat. Yet, the substantial dependence between predictors remains worrisome. There is ample reason to consider further the predictors that appear to matter most, which also tend to differentiate the actual June heat wave from the faux heat waves in June and July. Another data analysis approach may be needed.
7 Applying Genetic Algorithms To The Fitted AIRS Data
We have shown that the distribution of gain scores from the actual extreme heat wave and from Faux Heat Wave 1 differ in important ways. There also is some evidence that for the actual and faux heat waves, precursor variables 14 days earlier can differ as well. But dependence between the 14 day lagged predictors makes claims about such differences somewhat tenuous.
@@ 7 Applying Genetic Algorithms To The Fitted AIRS Data
The statistics literature has long suggested ways, inconclusively, to think about and analyze variables that are highly correlated (Alin, 2010). Some recent work seems especially relevant for our setting: variable selection combined machine learning (Chan et al., 2022). But all approaches we have seen to date accept the data on hand as the venue for progress. The results are a long way from closure.
@@
If there were such a thing as an ideal extreme heat wave and such a thing as an ideal set of warm summer days, perhaps the empirical messiness of heat wave drivers could be reduced substantially. Recall that the June PNW faux heat wave had many gain scores similar to those of the June actual heat wave. This is not surprising because surface temperatures are produced by normal weather and topographical forcings likely combined with dynamic forcings when a heat dome arrives. Ubiquitous factors such as earlier surface temperatures are still in play. Perhaps for similar reasons, many of the predictors available in the AIRS data must be highly correlated making it extremely difficult to isolate their separate effects. Might abstractions from extreme heat waves and usual summer days help? Under the right circumstances perhaps an ideal extreme heat wave might have only a few dominant predictors and an ideal garden-variety series of very warm days also might have only a few, but different, dominant predictors.
@@
We can borrow from the social sciences the concept of an “idea type” or a “pure type” (Garth and Mills, 1958) that abstracts the essential features of some phenomenon much like many key concepts in climate science. Then, perhaps a genetic algorithm can be employed to construct two ideal types: a synthetic population of extreme heat wave grid cells and a synthetic grid cell population of normal sequences of very warm summer days. Comparisons between the two might allow for a more definitive reading of their similarities and differences.

### v3

@@ 5 Temperature Changes from Pre-Test to Post-Test
We transition to empirical analyses by separately examining the PNW heat wave summer temperatures and a rendering of routine PNW summer temperatures. For each grid cell, the gain score is the post-test mean temperature minus the pre-test mean temperature. This is just a routine difference in means computed separately for the actual heat wave and for each faux heat wave. Despite the name, gain scores can be positive or negative and are often called “change scores;” a gain score is the change from pre-test to post-test. For the analyses immediately to follow, attention centers on the change between the 4-day pre-test mean surface temperature and the 4-day post-test mean surface temperature.
We stress that although causal processes are involved, no causal effects are being explicitly estimated from the gain scores (Tennant et al., 2022). We are capitalizing on pre-test/post-test associations to help simplify the analyses. For example, some grid cells are routinely warmer or cooler than others because of features such as urban heat islands, elevation, type of ground cover, and proximity to the Pacific Ocean or the Cascade Mountains. On moderate time scales, these location features do not change. Therefore, they cannot by themselves be directly associated with extreme heat waves that are necessarily dynamic over a period of several days; a constant cannot be associated with a variable. We treat them as fixed location features controlled by the use of gain scores (Maris, 1998, Maxwell et al., 2018). Only temporally dynamic variables can directly matter.
@@
5.1 Comparisons Between Gain Score Distributions for the Actual Heat Wave and Faux Heat Wave 1
@@
Research on extreme heat waves often starts with comparisons between surface temperature distributions and common theoretical densities such as the extreme value distribution (Touchette, 2009; McKinnon and Simpson, 2022; Zhang et al., 2024) Figure 3 shows the grid cell, surface temperature gain scores of means histogram for the extreme heat wave, for the Faux Heat Wave 1 and for their overlap. The gain scores condition directly on pre-test surface temperatures and, therefore indirectly on fixed temperature-related features of grid cells such as proximity to an urban heat island.444 Statistical interactions between dynamic processes and fixed features of a grid cell or the period (i.e., day) certainly can occur and are not controlled by the pre-test conditioning (Lui, 2016: section 1.4). Estimating such relationships can be very challenging and is beyond the goals of this analysis, necessarily constrained by what the AIRS data measures.
@@
Neither the actual heat wave histogram nor faux heat wave histogram has much resemblance to the canonical normal distribution. There is far too much mass toward the centers of both histograms. Nor is either distribution strongly skewed to the left or the right. There is nothing to suggest an extreme value distribution. The attempt to isolate the role of dynamic temperature process does not lead to popular “off-the shelf” distributions in gain scores.555 These are “eyeball” assessments that require some care because the aspect ratio of all three histograms produces rather flat-looking frequency profiles. When the histograms have an aspect ratio of 1 to 1 in height to width, the eyeball impression is close to a rectangle that is taller than it is wide. The 1 to 1 aspect ratio led to distorted displays when they were collected on a single page.
One might conclude that the two histograms are much alike. However, the extreme heat wave histogram has 3 of 120 grid cells with negative values and an overall gain score mean of 3.8. The faux heat wave distribution has 44 of 128 grid cells with negative values and an overall gain score mean of 1.2.666 There are less than 144 grid cells in either histogram because when a dataset for each was built with all of the predictors, there are 20 or more observations with missing data. Missing data can occur because of cloud cover or intermittent technical malfunctions. Missing observations in the AIRS data are an ongoing challenge. One can explore further how the distributions compare by plotting the two histograms in the same axes. The bottom display shows the distributional offset in more detail. The red bars are taken from the histogram for the extreme heat wave gain scores. The blue bars are taken from the histogram for the faux heat wave 1 gain scores. The purple bars show the overlap.
@@
In the interest of space, Faux Heat Wave 2 is not yet discussed. But its histogram looks a lot like the histogram for Faux Heat Wave 1. We make more use of Faux Heat Wave 2 later to replicate some important findings for Faux Heat Wave 1. The key point now is that when examining histograms, the role of a 4 day lag versus a 30 day lead does not seem to make much empirical difference in the gain score histograms.
6 Some Important Predictor Distribution Differences Between the Actual Heat Wave and the Faux Heat Wave 1
@@
One might wonder whether the actual heat wave and Faux Heat Wave 1 have dynamic precursors that differ as well. We focus on dynamic predictors because only they can forecast a heat wave, which are also dynamic. Such precursors were listed in Table 1. To proceed, however, we need first to consider a bit further how the rare events problem might be in part addressed.
@@
6.1 Mitigating The Rare Events Problem When Comparing Predictor Distributions
The relative infrequency of extreme heat waves depends on the population from which the data were realized. For example, if one considers all days from 2017 through 2021 in the Pacific Northwest as the set of days from which the data were realized, there would be one realized extreme heat wave.
@@ 6.1 Mitigating The Rare Events Problem When Comparing Predictor Distributions
Labeling all four of the last days in June of 2021 as heat wave days, there would be 4 extreme heat wave days among more than 1800 days overall. Based on the data, if one never predicted an extreme heat wave, that prediction would be correct for over 99.8% of the days without using any predictors at all. It would be, therefore, almost impossible to uncover the roles of predictors because fitting is already close to perfect. Even if the days from which the data were realized included all just 2021, a forecast of no heat wave would be correct for 98.9% of the days. The fitting problems are nearly as damaging.
@@
There is more promise considering only the 30 days in June of 2021. A no heat wave forecast will be right for “only” 86.7% of the days. This degree of imbalance might be statistically manageable. We can make further progress employing a within-subject research design (Maxwell et al., 2018: section III) that uses the same grid cells as the study units for the post-test, pre-test, and precursors measured two weeks earlier. Figure 3 avoids in a transparent (even trivial) manner the rare event problem within any of the three research design configurations because no comparisons between a true heat wave and a faux heat wave are made.
@@
There is also no evidence for rare event problem when making comparisons across the three design configurations. This will become important later, especially in Section 8. If the actual heat wave is compared empirically to either of the two faux heat waves, there are, in principle (i.e., without yet considering missing data), 144 true heat wave grid cells and 144 faux heat wave grid cells. If the extreme heat wave scenario is compared to a pooled version of the two faux heat waves, there remain 144 grid cells in the heat wave scenario and then, 288 grid cells in the combined faux heat wave scenarios. Because of the within-subject design, the PNW extreme heat wave in these data is not rare.
@@
There is a price to pay for these design remedies to the rare events problem. Perhaps most important, we use a form endogenous sampling when the three design configurations in Figure 2 are imposed to all grid cells. It is unlikely that any study of rare extreme heat waves could be undertaken with data from a naturally-occurring population in which the probability that a realized case would be found in any one of three different study coditions is exactly $1/3$. Forcing the data to be balanced in this manner can produce misleading statistical results because of a mismatch between joint distribution of the data and the joint probability distribution of the population. (Manski and Lerman, 1977; Manski and McFadden, 1981). We will return to this problem later to consider several possible solutions. For now, we continue to descriptively unpack how extreme heat wave precursor distributions can differ from faux heat wave precursors distributions. Ideally, such information subsequently will help when real forecasting is undertaken.
Figure 4: Comparisons Between Three Important Predictors For The Actual Heat Wave and Faux Heat Wave 1
6.2 Some Predictor Comparisons Between the Actual Heat Wave and the Faux Heat Wave 1 or Faux Heat Wave 2
Figure 4 compares predictor histograms from the actual heat wave and Faux Heat Wave 1. At the top is a comparison between distributions for tropopause height plotted red for the true heat wave, in blue for the Faux Heat Wave 1, and in purple for the overlap. The distributions differ in location, variability and shape. 14 days before both the actual heat wave or Faux heat wave 1, tropopause height was typically substantially higher in the actual heat wave cells. In the middle, there are also differences in distributional location, variability and shape. 14 days before both the actual heat wave or Faux heat wave 1, atmospheric temperature at level 8 was typically substantially higher in the actual heat wave cells. At the bottom, there are again distributional differences in location, variability and shape. 14 days before both the actual heat wave or Faux Heat wave 1, atmospheric water vapor at level 8 was typically substantially higher in the actual heat wave cells.
@@ 6.2 Some Predictor Comparisons Between the Actual Heat Wave and the Faux Heat Wave 1 or Faux Heat Wave 2
One repeated the exercise but replaced Faux Heat Wave 1 with Faux Heat Wave 2. The faux heat wave now is 30 days after the true heat wave. Consequently, the risks from carry-over contamination are dramatically reduced. On the average, however, weather conditions in June and July may differ. There can be less rain and lower soil moisture in July, for instance. According to NOAA National Centers of Environmental Information, 1.9 inches of rain fell in Seattle in June of 2021, and 0.0 inches of rain fell in Seattle in July of 2021 (https://www.ngdc.noaa.gov/).777 In the winter months, the rainfall in Seattle is much more substantial.
Despite such possibilities, the same comparisons produced similar results. Perhaps neither the difference in pre-test to post-test separation time nor the differences in weather patterns between June and July matter much. It also may not matter if the actual heat wave followed a faux heat wave or the reverse. No evidence of carry-over, order problems is found.
There probably are many other variables having distribution differences depending on the circumstances in which summer temperatures materialize. Nevertheless, the comparisons between histograms offer some additional signs that extreme summer heat waves might differ structurally from ordinary summer heat. Work on quasiresonant amplification is a provocative illustration (Petoukov et al., 2012; Mann et al., 2018; Li et al., 2024). However, the empirical details are hard to pin down because there is substantial dependence between many such predictors, especially the when same variable is measured at different altitudes. Proximate values can be highly correlated. It is also important to not uncritically accept any answer to the “compared to what” question. There are no doubt different kinds of “non-heat-waves” to which comparisons can be made. We picked comparisons by temporal and spatial proximity. There are perhaps better, or at least complementary, ways to proceed.
7 Applying Genetic Algorithms To The Gain Scores
@@
The statistics literature has long suggested, inconclusively, ways to think about and analyze variables that are highly correlated (Alin, 2010). Some recent work seems especially relevant for our setting: variable selection combined with machine learning (Chan et al., 2022). But all approaches we have seen to date rely on the data on hand as the venue for progress. The recommendations are a long way from closure.
@@
If there were an ideal extreme heat wave and an ideal set of ordinary hot summer days, perhaps the empirical messiness of heat wave drivers could be reduced substantially. Recall that the June PNW faux heat wave had many gain scores similar to those of the June actual heat wave. This is not surprising because surface temperatures are produced by normal weather and topographical factors likely combined with dynamic forcings when a heat dome arrives. Might abstractions from extreme heat waves and usual summer days help?
@@
We borrow from the social sciences the concept of an “idea type” or a “pure type” (Garth and Mills, 1958) that abstracts the essential features of some phenomenon much like many key concepts in climate science. Then, perhaps a genetic algorithm can be employed to construct two ideal types: a synthetic population of extreme heat wave grid cells and a synthetic population of normal warm summer day grid cells. Comparisons between the two might allow for a more definitive reading of their similarities and differences.
7.1 Constructing Algorithmic Survival Functions
Genetic algorithms generally require as input a survival function.888 “Survival functions” are also known as “fitness functions” or “selection functions.” To construct a survival function separately for the heat wave gain scores and for the gain scores from Faux Heat Wave 1, we considered two forms of widely used and successful supervised learning: random forests (Breiman, 2001a) and stochastic gradient boosting (Friedman, 2001). Both can be seen as variants of nonparametric regression, are very easy to train, and often the default tuning parameter settings work well. A complication is that there is strong dependence between many of the AIRS predictors. For example, the measured temperatures at different altitudes are often correlated at values above .95.
Random forests handles strongly correlated predictors relatively well because it is comprised of a large ensemble of regression or classification trees with the partitioning for each tree undertaken using a small random subsets of predictors. Sampling predictors (without replacement) for each split can dramatically reduce the impact of high multicollinearity.999 Random forests usually randomly samples cases when each tree is grown and predictors at each split in a tree. Random sampling of cases can be done with or without replacement. Random sampling predictors and cases is designed to increase the independence between trees. Greater independence facilitates the averaging over all trees in the forest. We recently learned that gbm3 in R now allows sampling predictors too, and along its R predecessor gbm also allows random sampling of cases. In R, the choice between random forests (i.e., randomForests) gradient boosting now depends largely on details that are peripheral to this discussion. We settled on the implementation of random forests in R, ported by Andy Law and Matthew Wiener (Liaw and Weiner, 2002) from the original Fortran by Breiman and Cutler (Breiman, 2001a), which includes several important auxiliary algorithms that can dramatically increase algorithmic transparency.101010 We briefly considered using deep neural networks through the keras3 package in R authored and maintained by Tomasz Kalinowski. Unlike random forests algorithm, neither the predictors nor the observations are randomly sampled. The vastly different units for some of the predictors caused problems for the back propagation. Dependencies between the predictors also caused computational difficulties. Once both problems were addressed, tuning was able to proceed. For all practical purposes, the fitting performance was about as good as random forests’. But, some of the auxiliary procedures available for random forests, which shine some light into the algorithmic black box (Molnar, 2022), were not readily available. For example, the widely applicable partial dependence plots procedure pdp in R, authored and maintained by Brandon Greenwell, apparently requires that a custom prediction function be coded when used on neural networks output.